# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 6 Summary
We could theoretically underpin (see Fig. 9) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[1]. However, not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation on the membrane, as well as polarization in the electrolyte surrounding the membrane. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Considering the presence of electrolytes on the two sides of the membrane, a correct value for the resting potential can be derived numerically from the first principles. Those layers have a decisive impact on the neuron’s electrical operation. The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [8] of science must be considered.
By introducing the theoretical idea that biological currents are slow, we underpinned the experimental observation by Hodgkin and Huxley [65] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". Recognizing explicitly that biological systems use slow currents of charged ions instead of electrons solved some old mysteries and defied old fallacies.
Appendix A Speculations
@@
we can estimate how the pressure of the neural cell increases due to the rush-in changes at the beginning of the AP. As evidence shows, the local potential at the internal surface of the membrane increases by $\Delta U=100\ [mV]$. This increase means a change in the force acting on an ion (see Eq. (48)) by $1.6*10^{-12}\ [N]$. When we assume $10^{7}$ rush-in ions and unchanged cell size, the total force acting on the membrane increases by $1.6*10^{-5}\ [N]$. This change in force means on the neuron’s $8*10^{-9}\ [m^{2}]$ surface (see [30], page 12) a pressure change
@@
A way to estimate the ions’ rush-in speed is that one assumes that the Stokes-Einstein speed describes the ions’ speed ([49], Fig. 3 quantitatively underpins the hypothesis) both in axons and ion channels. If so, the ratio of the potential gradients equals the ratio of the speeds of ions. One can estimate that the 100mV100~mV AP on the $50~\mu m$ AIS generates a $2*10^{4}~\frac{V}{m}$ electrical gradient and [49] measured $20~\frac{m}{s}$ for the speed of the AP. By assuming that the electrical field across the membrane is $10^{7}~\frac{V}{m}$ the speed at the exit of the ion channels can be estimated as $10^{4}~\frac{m}{s}$.
@@
We can also estimate the work done when the rush-in charge extends the neuron, as the $\Delta P\times\Delta V$. We consider that $\Delta P$ remains constant, and we calculate the change of the volume as neuron’s surface $10^{-8}\ m^{2}$ multiplied by the change of the radius $\Delta r=10^{-9}\ [m]$ multiplied by the pressure $\Delta P$, that gives $E_{mech}=2*10^{-14}\ [J]$. If one assumes that the pressure is of entirely mechanical origin [70], one must assume that the pressure is $10^{6}~\frac{N}{m^{2}}$. We can also estimate the energy $E_{electr}=\frac{1}{2}C*(\Delta V)^{2}=1.4*10^{-12}\ [J]$, see also [14]. To estimate the time of changing the pressure, one can assume that the collision force of a $Na^{+}$ ion provides the pressure (the shock wave) and it can be calculated as $F=m\frac{\Delta v}{\Delta t}$. Given that Eq. (48) provides the force and one must assume that the ion loses its $2*10^{5}~\frac{m}{s}$ speed, one arrives that it generates the shock wave in $\Delta t=5*10^{-8}~seconds$.
@@
The consequences of the changes in the electric charge are large enough to explain why pressure wave and other mechanical changes [37] also start at the beginning of the AP, and other (such as optical, density) changes are accompanied by it. Nature invests energy also in the conventional way, a term $V\Delta P$, into the thermodynamics of neural operation; not only in the form of storing energy in the changed electric field. Thermodynamics and electricity, not to mention elasticity, must not be separated when discussing the neuronal energy business.
@@
In line with the experimentally measured energy consumption and our calculation in section 4.5, the AP cycle consumes energy in the range of $10^{-7}\ [J]$. That is, practically all the energy is invested as elastic energy, but the modulus is exceptionally high. The primary reason for neuronal operation is the appearance of new ions in proximity to the membrane, but it creates a secondary, vast mechanical shock wave. By considering the electric repulsion force between the particles in the slow current, which is entirely omitted in the classic theory, we open the way for explaining the observed thermodynamic and mechanical changes. Instead of alternative disciplinary theories, cooperation between the classic disciplines is needed.
@@
For thermodynamic distributions, one can interpret the "temperature of individual particles" with mean kinetic energy $E$ as $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is, the temperature is directly proportional to the kinetic energy. The energy arrives at the neuron in the form of $10^{7}$ $Na^{+}$ ions, so the average energy a single ion consumes is $2.5*10^{-14}~[J]$. One can assume that most (99%) of that energy is consumed for moving the ion against the Stokes-Einstein force in the viscous medium, so after exiting the membrane, the speed of the ions is $\sqrt{\frac{2*2.5*10{-16}}{3.82*10^{-26}}}\approx 10^{5}~[m/s]$, which is well above the $\approx~500~m/s$ average speed. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electric potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature (the temperature is more that two orders of magnitude higher, but the proportion of those "hot" ions is about six orders of magnitude lower, so one can expect temperature change up to the range of fragments of millikelvins, depending on the measuring conditions). In the second phase of the AP, those more energetic ions that provide excess voltage (and so: excess local potential) above the resting potential leave the membrane’s proximity through the AIS, so the temperature decreases in the second phase to its original value; another effect that can contribute to the general heat production and adsorption discussed above. In the original publication about solitons (essentially pressure waves) [78], a lot of measured and theoretical data are discussed. They are in line with our conclusions, among others, $80\ \mu K$ measured temperature change; furthermore, that the electric energy is more than an order of magnitude smaller than the elastic energy (called soliton energy there).

### v3

@@ 8 Summary
We could theoretically underpin (see Fig. 9) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[17]. However, not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation at the membrane and to polarization in the electrolyte surrounding it. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Those layers have a decisive impact on the neuron’s electrical operation.
The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [31] of science must be considered. We successfully solved the mystery of the heat absorption of the neuron and derived a method for the thermodynamic description of the neuron (albeit based on parameters that can be measured electrically). Furthermore, its energy consumption and operational efficiency. We showed that, due to the disciplinary properties, one cannot derive the thermodynamic entropy using the usual methods of information processing.
Competing interests
The authors declare that they have no competing interests
Funding
The research did nor receive any support.
Appendix A Dichotomies in neuron modeling
Paper [9], which, in the context of this section, tends to be the precursor of our paper, sheds light on some remarkable dichotomies. Scrutinizing them can result in "a sound basis for unification of the physics of nerve impulses". By deeply agreeing with the fact that they have a "potential impact on our understanding of (the physical nature of) neuronal signaling", we list and uncover more dichotomies; furthermore, we uncover their fundamental, partly philosophical reasons.
A.1 Science methods
For centuries, science developed its methods for deriving abstract concepts by reducing the features of a real object to an abstract one that cannot be reduced further, such as mass and charge. It derived laws for the forces acting on those abstract objects, such as Newton’s universal law of gravitation and Coulomb’s law of electricity. Then one could apply Newton’s laws of motion. Experience shows that the generated forces, independently of their origin, can be summed, and one can apply the laws of motion by using the resultant force. Ions are special from the point of view of reducing the ’material point’ to one single abstraction: the ions are charge and mass simultaneously, without a further possibility of reduction. Notice that in the electrical abstraction, no mass is present, so one can use the equations assuming ’instant interaction’, which in biology lead to non-physical explanations of the observations (such as ’delayed current’). In the mechanical/thermodynamic abstraction, mass is to be moved, making the finite-speed interaction evident. The same physical phenomenon, the interaction (or movement) of ions is described using an ’infinite’ electrical speed and a million times less mass propagation speed, respectively, which leads to an unresolvable discrepancy, given that physics is not prepared to handle different speeds in the same interaction event [6]. This item is listed in Table 1 in [9] as the "Electrical vs Mechanical" dichotomy.
The low speed of ions in electrolytes introduces further problems. The farther parts of the cell will see any change in the local value of the state variables with a delay. The material transport represents simultaneously mass and charge, so the transport itself changes the gradients. This process keeps the entire volume of the electrolyte in (more or less) continuous change (which makes life). Furthermore, biological objects inside the cell can absorb ions and charge up. With their potential, they alter local gradients, accelerating or decelerating ions. Not to mention that biological objects can be active in the sense that (depending on the environmental conditions) they can let ions from one separated volume part into the other.
The closed volume of the biological objects is also a problem of finite resources. In biology, we cannot use the fundamental assumption of physics that, although a field acts on the ion in question, the ion does not affect the field (that the other ions generate). The transferred ions decrease the field in the volume they departed from and increase it in the volume where they arrived. Due to the field-dependent speed within the electrolyte, we must consider the autonomous change between the microvolumes where the ion traverses.
These processes are what E. Schrödinger coined as "the construction is different from anything we have yet tested in the physical laboratory[7]. Consequently, measurements must be designed and carried out with care; the routine methods used for measuring objects from inanimate nature, cannot surely be applied to living objects.
The interdependent behavior of charge and mass has an interesting mathematical consequence as well. The definition of a partial derivative of a function of several variables is its derivative with respect to one of those variables, with the others held constant. In the case of ions, changing the function with respect to mass or charge means simultaneously changing the other; that is, the partial derivatives depend on each other. In other words, one cannot calculate the partial derivatives; only the total derivative. As a consequence, equations that calculate the partial derivatives of concentration and electrical potential with respect to time are either incorrect or approximations. Classical mathematics cannot be applied. Consequently, in the case of ions and electrolytes, one must not use the well-established concepts (enthalpy, entropy, etc) of thermodynamics in unchanged form.
Although science is aware that the apparently continuous matter in nature cannot be divided infinitely, even though it knows that the abstracted discrete ’material points’ (as A. Einstein coined) have well-defined discrete values, it occurs rarely that both views must be applied in describing a single phenomenon. The continuous and discrete approaches (also called macroscopic and microscopic views) seem to be independent of each other. Connecting those views was one of the tasks performed by thermodynamics for particles with no long-range interactions. However, there is no similar discipline for electricity, although the behavior of charge carriers is similar to that of neutral discrete particles. It is not evident for electrons, but significant for ions. This item is listed in Table 1 in [9] as the "Macroscopic vs Microscopic" dichotomy.
Thermodynamics provides a framework for handling ions and determining their thermodynamic properties. However, as good thermodynamic textbooks (including the one on the thermodynamics of the membrane [84]) emphasize, thermodynamics derives its concepts for non-interacting particles, so one cannot expect its validity for ionic solutions [6] (Boltzmann assumed that, in the absence of long-range interaction between the particles, the sizes of cells in the phase space do not change). In addition, he required the presence of a vast number of particles in a homogeneous, isotropic, infinite volume, which is typically not the case in biology.
The dichotomy about reversibility is closely related to the exclusively electrical nature of the HH theory. In the framework of that theory, Ohmic currents flow through resistors that irreversibly dissipate heat due to friction, no matter in which direction ($W=I^{2}\times R$) the ion currents flow. By introducing the concept of ’delayed current’ and the hypothesis that some hidden power controls the operation of neurons by changing their conductance, physiology gave rise to the fallacy that science and life sciences are almost exclusive fields. Furthermore, their (unintended) model provokes questions (for a review see [12]) whether it is a model at all, and what controversies it delivers. As Hodgkin wrote [83]: "Hill and his colleagues found [22] that an initial phase [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." Since that invention, experimental evidence from the missing "leakage current" [45] to the conversion between elastic and kinetic energies [65] to demonstrating the pressure-wave-like behavior of AP witnesses that at least part of the phenomenon is of thermodynamic origin; that is, it requires a cross-disciplinary explanation. For discovering the reciprocal relations in thermodynamics [87], also in electrolytes, Lars Onsager was awarded the 1968 Nobel Prize in Chemistry. Those relations, together with the theoretical understanding, manifesting in the cross-disciplinary Nernst-Planck relation, paved the way to a combined theory. However, no thermodynamic explanation has been given so far. Our discussion provides, in line with the experimental results, a complex electrical/thermodynamic description of neuronal operation.
Those dichotomies are rooted in deeper layers of science. Classical physics is based on the Newtonian idea that space and time are absolute, so everything happens simultaneously. Consequently, when their objects interact, it must be instantaneous; in other words, their interaction speed is infinitely large. Furthermore, electromagnetic waves with the same high (logically, infinitely high) speed inform the observer. This self-consistent abstraction enables us to provide a "nice" mathematical description of nature in various phenomena: the classical science. In the first year of college, we learned that the idea resulted in "nice" reciprocal square dependencies, Kepler’s and Coulomb’s Laws. We discussed that the macroscopic phenomenon "current" is implemented at the microscopic level by transferring (in different forms) discrete charges; furthermore, that solids show a macroscopic behavior "resistance" against forwarding microscopic charges. We also learned that without charge (and, without atomic charge carriers), neither potential nor current exists. We did not learn, however, that thermodynamic forces can move the ions, given that they have inseparable mass.
Physics notoriously suffers from a lack of handling different simultaneous interactions; facing such a case leads to misunderstandings, debates, and causality problems. Such a famous case is the entanglement speed. In that time, E. Schrödinger introduced his famous law of motion in quantum mechanics entirely analogously to how I. Newton introduced his Laws of Motion. Similar to the Newtonian ’absolute time’, the quantum mechanical interaction is supposed to be ’instant’ (this is the price for having ’nice’ equations in classical and quantum mechanics), i.e., its speed is supposed to be infinitely high. However, at that time, it was already known that the speed of electrical interaction (propagation of electromagnetic waves) is finite. So if an object has quantum-mechanical interaction (aka entanglement) and electrical interaction simultaneously, the corresponding forces act simultaneously but reach the other object at different times. The entanglement arrives instantly; the electromagnetic effect arrives at a time we can calculate from the interaction speed and the spatial distance between the objects. This effect leads to causality problems: the two interactions of photons entangled earlier in an exploded supernova should be measured at two different times, meaning a "spooky remote interaction" as A. Einstein coined, and leads to contradictions such as the Einstein–Podolsky–Rosen paradox. Actually, the issue stems from the improper handling of mixing interaction speeds: the Schrödinger-equation introduces the infinitely large interaction speed, while the electromagnetic (EM) interaction has a finite speed.
The confusion and question marks in connection with describing life by science mostly arise from the interpretation of notion ’speed’ in physics. When discussing the underlying physical laws for biology, we go back to the fundamental physical concepts instead of taking over the approximations and abstractions used in the classical physics for non-biological matter and less complex interactions. As we emphasized many times, we construct laws and conclusions based on simplified abstractions about nature in all fields of science. The notions and laws depend on the circle of phenomena we know and want to describe. The Newtonian and Einsteinian worlds are basically distinguished by considering speed dependence that actually means explicit time dependence. Interesting consequences are that in the Einsteinian world, the mass is not constant, time and space are not absolute, and so on. We can be prepared for some similar counter-intuitive experiences in physiology: "we must be prepared to find it working in a manner that cannot be reduced to the ordinary laws of physics"[7]. Here, we scrutinize the fundamental concepts and discover differences between physics and biology as consequences of the required different abstractions and approximations.
During our college studies, we mentioned that light is an electromagnetic wave with a vast but finite propagation speed. Still, we forgot to highlight that, at the same time, it is the propagation speed of the electric (, and magnetic, and gravitational) interaction force fields as well. The effect of "Retarded-Time Potential" is also known in physics and communication engineering. Algorithms "marching-on-in time" and "Analytical Retarded-Time Potential Expressions" are derived to handle the problem; for a discussion, see [88]. Telegrapher’s equations (unfortunately, also used to describe biological signal transfer) explicitly assume a finite propagation speed, millions of times slower than the (implicitly assumed) EM interaction’s. The issue is not confined to large distances: designers of micro-electronic devices also must consider the effect: they introduced clock time domains and clock distribution trees; see, for example [89, 90].
Science uses ’instant’ in the sense that one interaction is much faster than the process under study; we consider the faster interaction as instant. The approach of classical science is based on the oversimplified approximation that the interaction speed is always much higher than the speed of changes it causes and that the processes can always be described by a single stage. In our approach, for biology, we put together a series of stages to describe the observed complex phenomena, where the stages provide input and output for each other, involve more than one interaction speed, and use per-stage-valid approximations. We simplify the approximations by omitting the less significant interactions and introduce ideas for accounting for the different interaction speeds. This way, we reduce the problem to a case that science can describe mathematically. This procedure is fundamentally different from applying some mathematical equations derived for an abstracted case of science to a complex biological phenomenon without validating that we use the appropriate formalism.
A.2 Complete measurement
In physics, measurement means the determination of something relevant to the process under study. A ’complete’ measurement measures all relevant quantities. The different disciplines of physics restrict the measured quantities to the ones which are, in general, that the discipline studies. The remaining quantities remain outside the scope of the discipline. The fundamental physical quantities of mechanics are length, mass, and time, which form the basis for defining all other quantities like velocity, force, and energy in that field. The fundamental physical quantities in thermodynamics are Temperature, Energy (Internal Energy, Heat, Work), and Entropy, which characterize systems at equilibrium and describe energy transformations; Pressure and Volume also being key variables. Those of electricity include electric charge (Coulomb), the basis of all electric phenomena; electric current (Ampere), the rate of charge flow; voltage (Volt), the potential difference driving current; resistance (Ohm), opposition to current; power (Watt), the rate of energy transfer; and energy (Joule), the capacity to do work. As seen, there is little overlap of the studied quantities. A disciplinary study (such as the electrical and thermodynamic ones in physiology) provides an incomplete set of measured quantities. None of the disciplines alone can describe electrolytes given that they consider a different and incomplete set of physical quantities. As discussed in connection with the Onsager relations, a disciplinary (incomplete) measurement does not discover that an unexpected change (a miracle) in the value of one of the quantities belonging to another discipline is accompanied by a change in the value of the studied quantity. This change is clearly the case with the measurements performed in the spirit of HH: the values the electrical instruments provide are accompanied by the values of mechanical/thermodynamic quantities which are not measured, partly due to the obvious measuring difficulties and because they are outside the scope of the discipline. One cannot reduce ion-related phenomena to a single abstraction as thermodynamics does with mass or electricity does with charge. If one attempts (see the mentioned theoretical descriptions), one experiences that some quantity (the charge or the mass) changes in a uncontrolled way. As Feynman R. P [91] told, "the separation of fields, as we have emphasized, is merely a human convenience, and an unnatural thing. Nature is not interested in our separations, and many of the interesting phenomena bridge the gaps between fields". The disciplinary separation in classical science does not apply to the study of life.
A.3 Physics and biology
A common fallacy in biology is that physics cannot underpin the operation of living matter, citing E. Schrödinger. However, the claim falsifies his opinion by omitting the most essential word, ’ordinary’. Schrödinger wanted to emphasize the opposite: there is no new force (no unknown new interaction, as biophysics attempts to introduce "protein mechanism"), only that studying living matter requires different testing methods (and we add: different uses of physics’ concepts) in the physical laboratory. He suggested answering the question "Is life based on the laws of physics?” affirmatively, but expected to discover the appropriate forms of physical laws describing the ’non-ordinary’ (in our reading: non-disciplinary) behavior of living matter. No doubt, the basic concepts and terms must be interpreted precisely for living matter, much beyond the level we used to at the college level. However, after that reinterpretation, we can interpret features of living matter, although we need a more careful, cross-disciplinary analysis to do so. We must use appropriate abstractions and approximations for the phenomena, depending on the level required for the given cooperation of objects and interactions. Furthermore, we discuss some of the relevant terms and notions of physics, differentiating which approximation is appropriate only for physics (mainly electricity), and, instead, which approximation should be used for biology. As we discuss, biophysics translated the corresponding major terminus technicus words from the theory and practice of physics’ major disciplines, mainly from electricity, which were worked out for homogeneous, isotropic, structureless metals, and for strictly pair-wise interactions with a single (actually, ’instant’) interaction speed; to the structured, non-homogeneous, non-isotropic, material mixtures and for multiple interaction speeds. Those notions do not always have unchanged meaning, and how much they do, depends on the actual conditions. The precise meaning needs a case-by-case analysis.
The physical models consider infinitely large volumes, surfaces, distances; furthermore, and most importantly, instant interactions. Is the cell large enough to consider it infinitely large (at least on the scale using ions’ size); that is, to apply laws of science derived for the abstraction ’infinitely large’? When working with charge, we know that charge is quantized, while the macroscopic quantities voltage and current are continuous (derivable). Do cells contain a sufficient number of charge carriers to apply macroscopic notions? Do the thousands of times smaller ion channels transfer enough charge to speak about a well-defined current? When a couple of ions are transferred through an ion channel, do they significantly change the potential that accelerates them?
Science could serve as a firm base for all its disciplines. As we discuss, its disciplines use abstractions based on limited-validity approximations based on the same first principles. However, the approximations differ between biology and physics. In physics, some processes we observe are fast enough so that we can use the abstraction that they are essentially state transitions. In some cases, the approach can be –more or less– successful. For the slower, well-observable processes, we have the laws of motion that describe how processes occur under the influence of a driving force. We also found that nature is not necessarily linear (in the sense that it depends only on the mentioned quantities, not on their derivatives), which we can describe with "nice" mathematical formulas. A century ago, A. Einstein discovered that the approximations I. Newton introduced two centuries earlier are not sufficiently accurate for describing the movement of bodies at high speeds. In other words, a new paradigm, the constancy of the speed of light, must have been introduced that caused a revolution in physics and led to the birth of "modern physics".
Life, including the brain’s operation, is dynamic. As Schrödinger formulated, the "construction of living matter" differs from the one science used to test in its labs. The scientific abstraction based on "states" (i.e., on instant changes) fails for the case of biology, where "processes" happen (i.e., the changes are obviously much slower). The commonly used measuring methods, such as clamping, patching, and freezing, reduce the life to states. The related theories describe states with perturbation [92]. On the one hand, this technology fixes the cell in a well-defined, static state, enabling us to observe a static anatomic picture of the cell. On the other hand, it eliminates the dynamic processes from the theory, i.e., hides forever the essence of the life that the cell exists in a continuous change governed by laws of motion. Those methods stop the processes under study, thereby depriving them of their dynamicity and preventing measurement. It was forgotten that using feedback for stabilizing an autonomously working electrical system means introducing foreign currents, and this way falsifying its operation.
Furthermore, it is hazardous to introduce technically (and incorrectly) derived and misinterpreted macroscopic features and interpret them as fundamental electric notions. In general, instead of understanding and developing the proper scientific basis for the operation, they say that science cannot describe it. The idea of conductance has been introduced to neurophysiology almost a century ago. It was taken from physics, where the notion was derived for metals (conductors, instead of electrolytes). Since then, its original interpretation has been forgotten, and today (in contrast with physics), it has become a primary entity for describing electrical characteristics of biological cells. We explain how the right physics background enables us to discover wrong physical models and misinterpreted notions of physics in neurophysiology. Furthermore, the proper interpretation opens the way to the correct interpretation of neuronal information. We set up an abstract electrical/thermodynamic model of neuronal operation.
We derived the needed ’non-ordinary laws’ [6], which are derived by using the same first principles as the ’ordinary laws’, but are abstracted for the approximations valid for living matter. As we discuss, those ’ordinary’ laws were derived for strictly pair-wise interactions at very high speeds. In biology, we can observe interactions at a million times lower speed, in inhomogeneous, non-isotropic, structured material. Biology does not have the conditions for which physics derived its ordinary laws. By using the appropriate approximations for the biological cases, we can derive the required ’non-ordinary’ laws of physics, which laws are more complex to derive and use; furthermore, we must use several stages (with the approximations changing from stage to stage) instead of one single stage, as in the case of the ’ordinary’ laws. However, all laws follow the same principles.
Biology, and predominantly neuronal operation, produces examples where using wrong omissions in complex processes results in absolutely wrong results. In those cases, some initial resemblance between our theoretical predictions and our phenomena exists. However, the success in simple cases provides no guarantee that the model was appropriate: "the success of the equations is no evidence in favour of the mechanism" [14]. Finally, all laws are approximations, and the accuracy of verifying their predictions is limited. Several theories can describe the same phenomenon with the required accuracy. We also show in the section about the finite interaction speeds that the most well-known laws (from Newton, Coulomb, Kirchoff, etc.) are also approximations. They have their range of validity, although it is often forgotten.
One such neuralgic point of omissions and approximations is the vastly different interaction speeds. Furthermore, where the speed is considered at all, the same speed is assumed for all interactions. The laws are abstract also in the sense that, say, the objects in the laws of physics have either mass or electric charge, but not both. It is the researcher’s task to decide which combination of laws can be applied to the given condition. For example, one can assume in most cases that the speeds sum up linearly, except at very high speeds. Biology provides excellent case studies where different interactions shape the phenomenon, and special care must be exercised. We give a short review of history and kinds of interaction speeds.
Another point is that science started with the assumption that the non-living matter is continuous, although it was early discovered that there are the smallest pieces of that matter. When we reached that size, we experienced that different subsets of science laws describe that matter and the atoms they contain. It is one of the most challenging tasks to establish relations between those subsets. Again, we used abstractions that the continuous matter is infinitely large and that the isolated atoms are infinitely far from each other and from the external world. We also experienced the semi-infinite cases, and studied the behavior of surfaces and interfaces, which, again, is different from both that of the atoms and their large masses. Given that biological objects span the microscopic to macroscopic size range and are surrounded by surfaces, we must be prepared for the fact that no simple rules describe their behavior.
Neuronal operation is at the boundary, where sometimes, in the same phenomenon, one interaction can be interpreted at the macroscopic level, another must already be interpreted at the microscopic level. Furthermore, a series of stages (instead of a single state) and processes (instead of stages) describes the subject under study. Given the vital role of charge and current in neuronal operation, we provide their precise interpretations. Furthermore, we must consider that the processes happen in a finite volume, "within the spatial boundary".
Appendix B Speculations
@@
we can estimate how the pressure of the neural cell increases due to the rush-in changes at the beginning of the AP. As evidence shows, the local potential at the internal surface of the membrane is in the range of $100\text{\,}\mathrm{mV}$ in the resting state and increases by $\Delta U=$ $100\text{\,}\mathrm{mV}$ in the transition state. This increase means a change in the force acting on an ion (see Eq. (52)) by $1.6*10^{-12}\ [N]$. When we assume $10^{7}$ rush-in ions and unchanged cell size, the total force acting on the membrane increases by $1.6*10^{-5}\ [N]$. This change in force means on the neuron’s $8*10^{-9}\ [m^{2}]$ surface (see [40], page 12) a pressure change
@@
A way to estimate the ions’ rush-in speed is that one assumes that the Stokes-Einstein speed describes the ions’ speed (Fig. 3 in [14], quantitatively underpins the hypothesis, see [94]) both in axons and ion channels. If so, the ratio of the potential gradients equals the ratio of the speeds of ions. One can estimate that the 100mV100~mV AP on the $50~\mu m$ AIS generates a $2*10^{4}~\frac{V}{m}$ electrical gradient and [14] measured $20~\frac{m}{s}$ for the speed of the AP. By assuming that the electrical field across the membrane is $10^{7}~\frac{V}{m}$, one can estimate the speed at the exit of the ion channels as $10^{4}~\frac{m}{s}$.
@@
We can also estimate the work done when the rush-in charge extends the neuron, as the $\Delta P\times\Delta V$. We consider that $\Delta P$ remains constant, and we calculate the change of the volume as neuron’s surface $10^{-8}\ m^{2}$ multiplied by the change of the radius $\Delta r=10^{-9}\ [m]$ multiplied by the pressure $\Delta P$, that gives $E_{mech}=2*10^{-14}\ [J]$. If one assumes that the pressure is of entirely mechanical origin [80], one must assume that the pressure is $10^{6}~\frac{N}{m^{2}}$, which is well above the measured value. We can also estimate the electrical energy $E_{electr}=\frac{1}{2}C*(\Delta V)^{2}=1.4*10^{-12}\ [J]$, see also [45]. To estimate the time required to change the pressure, one can assume that the collision force of a $Na^{+}$ ion provides the pressure (the shock wave), and this can be calculated as $F=m\frac{\Delta v}{\Delta t}$. Given that Eq. (52) provides the force, and one must assume that the ion loses its $10^{4}~to~10^{5}\frac{m}{s}$ speed, one concludes that it generates a shock wave in $\Delta t=10^{-8}~to~10^{-7}~seconds$.
@@
The consequences of the changes in the electrical charge are large enough to explain why pressure wave and other mechanical changes [65] also start at the beginning of the AP, and other (such as optical, density) changes are accompanied by it. Nature invests energy also in the conventional way, a term $V\Delta P$, into the thermodynamics of neural operation; not only in the form of storing energy in the changed electrical field. Thermodynamics and electricity, not to mention elasticity, must not be separated when discussing the neuronal energy business.
@@
In line with the experimentally measured energy consumption and our calculation in section 7, the AP cycle consumes energy in the range of $10^{-7}\ [J]$. The five orders of magnitude between the total and the electrical energy seems to underpin that in addition to the $10^{-3}$ concentration of ions, only a small fraction of the ions (the ones in the layers near the membrane) in the volume participate in the electric activity of issuing an AP, while the entire volume in the mechanical activity. That is, practically all the energy is invested as elastic energy, but the elasticity modulus of the membrane is exceptionally high. The primary reason for neuronal operation is the appearance of new ions in the vicinity of the membrane, but this creates a secondary, vast mechanical shock wave. By considering the electrical repulsion force between particles in a slow current, which is entirely omitted in the classic theory, we open the way for explaining the observed thermodynamic and mechanical changes. Instead of alternative disciplinary theories, cooperation between the classic disciplines is needed.
@@
For thermodynamic distributions, one can interpret the "temperature of individual particles" with mean kinetic energy $E$ as $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is, the temperature is directly proportional to the average kinetic energy. The energy arrives at the neuron in the form of $10^{7}$ $Na^{+}$ ions, so the average energy a single ion consumes is $2.5*10^{-14}~[J]$. One can assume that most (99%) of that energy is consumed for moving the ion against the Stokes-Einstein force in the viscous medium, so after exiting the membrane, the speed of the ions is $\sqrt{\frac{2*2.5*10{-16}}{3.82*10^{-26}}}\approx 10^{5}~[m/s]$, which is well above the $\approx~500~m/s$ average speed. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electrical potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature (the temperature is more than two orders of magnitude higher, but the proportion of those "hot" ions is about six orders of magnitude lower, so one can expect temperature change up to the range of fragments of up to millikelvin, depending on the measuring conditions). In the second phase of the AP, those more energetic ions that provide excess voltage (and so: excess local potential) above the resting potential leave the membrane’s proximity through the AIS, so the temperature decreases in the second phase to its original value; another effect that can contribute to the general heat production and adsorption discussed above. In the original publication on solitons (essentially pressure waves) [15], a wide range of measured and theoretical data is discussed. They are in line with our conclusions, among others, $80\ \mu K$ measured temperature change. Furthermore, the electrical energy is more than an order of magnitude smaller than the elastic energy (called soliton energy there).